Let a = (2,-2, -1), b = (1, -2,2), c = (2,1,2). Find the component of a x b in the direction of c. 07 oj 00 O 21 O2
Answer #1

a=(2,-2,-1) b = (1, 2, 2) c = (2,1,2) = l ax ê Now J -2 K -1 2 2 1 -2 ai (-4-2) -] ( 4+1) tiê (- 4+2). axb =-bê-s9 - 2o - laxcomponent of a to along et in ax5). ł = (-6€ - 59 az ê ) . (2575 + 22) 14 3 3